Circular No. 7387

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE B. E. Schaefer, Yale University, on behalf of the QUEST collaboration

(see Schaefer et al. 1999, Ap.J. 524, L103), reports on 22 apparent supernovae discovered with the QUEST 16-CCD array camera on the Centro de Investigaciones de Astronomia 1-m Schmidt telescope at Llano del Hato, Venezuela. The total area searched was 264 deg^{2} between Mar. 1 and 15 to R = 20.8, with each clear night receiving independent drift scans through B, V, and R filters, so that the 22 objects all have well-sampled 3-color light curves over roughly two weeks. Each apparent supernova was found by subtraction of QUEST reference images taken on five nights between Feb. 8 and 14 and each new object has been confirmed via images taken on four or more nights in March. Five of the objects have no apparent host galaxy, to R ∼ 21.5. SNe 2000al and 2000aq peaked at R = 20.0 about Mar. 15; SN 2000am peaked at R = 19.6 around Mar. 14.

SN  2000 UT  α_{2000}  δ_{2000}  R  Offset
2000Y Mar. 4  7^{h}48^{m}13.^{s}00 − 1^{o}28′37′′.7 18.8 1′′ W, 8′′ N
2000Z Mar. 4 7 50 01.11 − 0 34 58.1 20.3 2′′ E, 10′′ N
2000aa Mar. 4 8 12 51.83 − 0 54 08.6 20.0  —
2000ab Mar. 4 8 27 05.82 − 0 15 26.9 20.2 2′′ W, 1′′ N
2000ac Mar. 3 8 58 22.78 − 1 40 04.9 20.6 1′′ E, 0′′ N
2000ad Mar. 3 9 14 00.30 − 2 16 33.7 19.9 1′′ W, 2′′ N
2000ae Mar. 9 11 45 11.40 − 1 38 11.2 19.5 1′′ E, 0′′ N
2000af Mar. 1 11 48 55.71 − 0 58 36.2 19.8 0′′ E, 0′′ N
2000ag Mar. 1 11 52 04.51 − 0 49 10.9 19.8 2′′ E, 1′′ S
2000ah Mar. 9 11 53 27.24 − 0 59 01.5 20.9 4′′ W, 6′′ N
2000ai Mar. 1 12 06 15.84 − 1 05 38.7 20.0  —
2000aj Mar. 9 12 37 01.22 − 1 49 46.7 20.3 2′′ E, 4′′ S
2000ak Mar. 1 13 06 28.73 − 0 13 30.0 20.7  —
2000al Mar. 1 13 11 49.65 − 1 21 19.4 20.6 1′′ W, 6′′ S
2000am Mar. 3 13 32 00.98 − 2 11 38.1 20.1 3′′ W, 2′′ S
2000an Mar. 10 13 46 23.94 − 0 06 34.2 20.3 1′′ W, 3′′ S
2000ao Mar. 3 13 57 49.80 − 0 01 04.8 20.4 1′′ E, 1′′ S
2000ap Mar. 3 14 09 23.65 − 0 35 01.7 20.0  —
2000aq Mar. 9 14 10 24.04 − 0 18 06.9 20.7  —
2000ar Mar. 1 14 12 11.89 − 0 29 05.2 20.3 1′′ W, 0′′ N
2000as Mar. 7 14 36 09.13 − 0 25 36.3 20.1 1′′ E, 2′′ N
2000at Mar. 9 14 55 31.00 − 1 53 11.2 20.6 4′′ E, 5′′ N

2000 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green